# Flexible-length Text Infilling for Discrete Diffusion Models

### v1

@@ 4.1 Continuous Position Diffusion
${\mathcal{L}_{pos}(\theta)=\mathbb{E}_{(z,t)}\left[Q_{t}(x_{t},y)\|v_{\theta}(z_{t},t)-(z_{0}-z_{T})\|^{2}\right]}$ (6)
We investigate two methods of initializing $z_{T}$ to begin generation. The first method is to sample $z_{T}\sim\mathcal{U}(-1,1)^{L}$, which we refer to as DDOT-random (DDOT-R). We then randomly select $l^{p}$ tokens from $z_{T}$ to serve as the positions for the prompt tokens. However, some areas of high density arise when randomly sampling which tend to map to pad tokens because there are less tokens in the corresponding ground truth region. Therefore, we propose DDOT-uniform (DDOT-U), which uniformly spaces out the prompt and infilled positions: $z^{p}_{T}=\text{linspace}(-1,1,l^{p})$ and $z^{r}_{T}=\text{linspace}(-1,1,l^{r})$.
During the sampling process, the token paths tend to follow straight trajectories because the ground truth velocity, $z_{0}-z_{T}$, remains constant over time. Therefore, we find it sufficient to use Euler steps during sampling to predict the positions. Combined with $\tau$-leaping Lou et al. , the straight position paths allow for favorable compute-accuracy tradeoffs and result in fast sampling. The total loss $\mathcal{L}$ is a simple linear combination $\mathcal{L}=\mathcal{L}_{tok}+\lambda\mathcal{L}_{pos}$.
@@ 4.2 Simultaneous Text & Position Diffusion
DDOT performs discrete text and continuous position diffusion simultaneously, as these processes operate independently in continuous time. We therefore predict both token value scores and position velocities in a single forward pass. This independence also enables simulation-free training by independently sampling token and position states at arbitrary timesteps. We summarize the training procedure in Appendix A.5
@@ 4.3 Implementation Details
Due to compute constraints, we implement our framework on SEDD Lou et al. . Our flexible position diffusion component is orthogonal to existing text diffusion methods and minimally alters the permutation of text tokens for infilling. This design allows DDOT to integrate seamlessly with various pretrained text diffusion models.
We extend SEDD, which is based on the Diffusion Transformer architecture Peebles and Xie (2023), with two additional modules. First, we introduce a learnable type embedding applied directly after the token embedding lookup. This embedding indicates whether a token is part of the prompt or the masked response ($x\in x^{p}$ or $x\in x^{r}$), which is critical for assigning each token to the correct OT flow. Second, we add a linear head at the end of the diffusion transformer to compute $v_{\theta}(z_{t},t)$.
To incorporate continuous positional information, we scale $z_{t}$ from the range $[-1,1]$ to match the context length of the original pretrained model (1024). We then use Rotary Position Embeddings Su et al. (2024), a standard technique in discrete diffusion models. Implementation details can be found in subsection A.3.
@@ 5.1 Experimental Setup
Datasets    We evaluate our approach on the One-Billion-Words and Yelp, following the preprocessing steps outlined in prior works on infilling and lexicographically constrained generation Miao et al. (2019); Zhang et al. (2020); Iso (2024). These datasets consist of examples with $1$ to $6$ keywords that must be infilled while maintaining their relative order to generate coherent sentences. In addition to randomly masking positions, we also introduce the "block" masking method that masks a single continuous chunk of text 00 to $L/2$ tokens long (32 for one-billion-word and yelp and 512 for CodeParrot). Finally, we apply the aforementioned masking methods to the Python subset of the CodeParrot dataset. Table 1 illustrates examples of this lexicographically constrained generation task.
@@
Model One-Billion-Word-Block Yelp-Block CodeParrot-Block CodeParrot-Random
@@
Model One-Billion-Word-Random Yelp-Random
@@
Training Details    To align the position prediction modules, we first finetune SEDD with the added modules on FineWeb-Edu Penedo et al. (2024). Afterwards, we further finetune on One-Billion-Word and Yelp datasests. For simplicity, we always keep all parameters unfrozen and simultaneously optimize both $\mathcal{L}_{tok}$ and $\mathcal{L}_{pos}$.
@@
In line with SEDD, we train our model in two configurations: small (90M non-embedding parameters) and medium (320M non-embedding parameters). DDOT-medium is on the same scale as CBART (406M parameters), and AutoTempalate-base (220M parameters). Following SEDD, we use the AdamW optimizer with a learning rate of 3e-5. We set $\lambda=10$ when using scaling. For each experiment, we either use 48 L40s (48 GB) GPUs, 80 A30 (24GB) GPUs, or 8 A100 (80GB) GPUs.
@@
Baselines    We compare our method against strong autoregressive (AR) and non-autoregressive (NAR) baselines. AutoTemplate Iso (2024), the state-of-the-art AR model, leverages the T5 Raffel et al. (2020) family of pretrained models. Specifically, AutoTemplate parses the lexicographically constrained generation task into a template that is autoregressively generated from left to right. The previous state-of-the-art NAR method, CBART He (2021), is built upon the BART Lewis et al. (2020) pretrained framework and iteratively inserts tokens into a sequence.
@@
We also introduce two diffusion-based models that follow the same training procedure as DDOT. Left Context (LC) concatenates all the prompt tokens to the left of the sequence and generates the response to the right of a separator token. Position Prediction (PoP) uses a SEDD model with a linear head that first predicts the positions of every token. Then, this sequence is fed through a finetuned fixed-position SEDD.
@@
Distribution Annealing    Many lexiocagphically constrained generation baselines including AutoTemplate and CBART use distribution annealing methods such as top-p, top-k, greedy sampling, and beam search. To provide a parallel to greedy decoding which always takes the top token probability, we anneal the distributions of our tokens values during sampling to only include the most probable token. Specifically, given the predicted probability of a certain token being mask $\hat{p}(x^{mask})$, we assign $1-\hat{p}(x^{mask})$ to the token value with the highest probability excluding the mask token. The rest of the token probabilities are set to 00. Greedy decoding in prior models (such as autoregressive) is deterministic, collapsing the tree of all generation paths into a single path. However, our annealing process maintains generation diversity (A.2) because the model must still sample from the annealed distribution with the top token value and the mask token. Whenever possible, we evaluate against the greedy decoding baseline.
Metrics  Following prior works Miao et al. (2019); Zhang et al. (2020); Iso (2024), we evaluate on BLEU-2/4 Papineni et al. (2002), NIST-2/4 Doddington (2002), and METEOR-v1.5 Denkowski and Lavie (2014), and success rate (SR).
@@ 5.2 Main Results
We present lexicographically constrained generation results in Table 2. Our approach uses greedy annealing and is compared against greedy decoding wherever applicable, including the CBART greedy decoding baseline. Our method achieves competitive performance with previous NAR models, approaching AR performance. Notably, our model achieves state of the art performance on most metrics among the diffusion baselines. Our method does well on block infilling, which may be more useful in real-world applications. Furthermore, we notice that DDOT scales well to longer sequences. Specifically, it frequently generates valid responses that include all the prompt words in the same relative ordering, as shown in the success rates (SR). In contrast, diffusion baselines quickly generate invalid responses as the number of prompt tokens increase (Figure 3, Table 2). However, in benchmarks with 6 or less prompt tokens, diffusion baselines maintain high SR (Table 3). This may be because the fixed-position models have room to correct generation when the ratio of prompt to response tokens is low.
Table 3 Compares results with previous works in the lexicographically constrained generation line of work. Since pretrained diffusion models lag behind AR models, an issue not unique to DDOT, we elect to focus on NAR models. DDOT performs on par with the previous SOTA models, and achieves higher SR than all diffusion baselines. Although DDOT underperforms LC and PoP in some metrics, we argue that one-billion-word-random and yelp-random over-index on the unrealistic task of generating text with only $1-6$ randomly spaced tokens, and that Figure 3 shows the broader trend when DDOT is scaled to more prompt tokens.
@@ 5.3 Analysis
In this section we investigate the effect of random versus uniform position initialization, the inclusion of OT coupling, and the impact of varying the number of sampling steps.
Figure 4: (a) and (b) exhibit Ground Truth Token Velocities. The velocities without OT (a) have a lot of crossing lines demonstrating instability in matching whereas the velocities with OT coupling (b) show almost straighter lines throughout the denoising process. (c) Performance tends to increase with sampling steps.
Position Initialization    In Table 2 and Table 3 we also explore the difference between DDOT-R and DDOT-U. In DDOT-R, pad tokens tend to cluster in areas of high density because the OT finds no match for them. However, the pad tokens in DDOT-U tend to be evenly spaced out. We find that DDOT-U consistently outperforms DDOT-R.
OT Coupling    To demonstrate the importance of OT coupling between source and target positions, we retrain the small version of DDOT without OT coupling and provide a quantitative comparison in Table 4. Models trained with OT coupling consistently outperform those using independent (random) coupling. We theorize that the OT coupling provides more signal about the ordering of tokens throughout the diffusion process. Specifically, the DDOT guarantees that the relative ordering of the prompt and generated tokens at any timestep is the same as the original ordering. In contrast, independent coupling requires the model to infer the original ordering of tokens—a challenging task given the numerous plausible orderings that can result from interspersed prompt tokens.
@@
Position Over Time    We qualitatively compare the ground truth token paths during training Figure 4. With OT coupling, token trajectories exhibit significantly fewer crossings, maintaining relative order throughout the generation process. In contrast, the independent coupling frequently permutes tokens. Visualizations of token paths during inference are available in subsection A.4
Number of Sampling Steps    One advantage of diffusion models over autoregressive models is their ability to exchange compute for accuracy by varying the number of inference steps. Figure 4 shows how the number of sampling steps influences lexicographically constrained generation performance. As the number of sampling steps increases, performance also increases.
Wall Time Analysis    We evaluate the inference speed of DDOT against the diffusion baselines on the One-Billion-Word dataset. Table 5 presents the wall-clock time per batch alongside BLEU-2 and BLEU-4 scores for an increasing number of sampling steps.
DDOT demonstrates significantly better efficiency. For any given number of sampling steps, DDOT is not only faster than LC and competitive with PoP in terms of raw speed, but also achieves substantially higher BLEU scores. Notably, LC must regenerate prompt tokens and therefore requires up to double the input sequence length. PoP also requires an additional forward pass to predict initial positions.
@@
The added modules to enable position prediction are lightweight, consisting of a linear head and two type embeddings. On the other hand, the LC baseline requires double the context length of DDOT because it must regenerate prompt tokens.
The OT calculation is highly efficient, taking 16 minutes and 11 seconds on an Intel Xeon 8462Y+ 64 core processor for the 10 billion token subset of FineWeb EDU. In practice, we stream the dataset, caching several OT couplings in advance without needing to preprocess the OT. With caching, it takes 4 minutes and 30 seconds to do 1000 training steps on an L40 gpu with a batch size of 256. Without caching, it takes 4 minutes and 27 seconds, a negligible difference.
@@ 6 Conclusion
In this work we introduce DDOT, the first discrete diffusion model capable of flexible-length text infilling by jointly denoising token values and positions. By incorporating optimal transport coupling, DDOT preserves token order while enabling dynamic position adjustment, addressing limitations in existing text diffusion models. Our experiments show that DDOT outperforms diffusion baselines and matches SOTA non-autoregressive models.
7 Limitations
DDOT inherits many drawbacks from previous discrete text diffusion methods Lou et al. ; Austin et al. ; Sahoo et al. (2024). First, diffusion backbones still lag behind autoregressive backbones. Since diffusion for text is an emerging field, this performance gap is expected. Similarly, the finetuned versions of these models (such as DDOT vs AutoTemplate) also exibit the same performance gap. Second, unlike autoregressive models, existing diffusion models cannot use a kv-cache to store previous activations. DDOT also inherits risks common between large-scale text models.
8 Ethics Statement
Artifacts  The artifacts that we used had public use licenses. Furthermore, we plan to release our code artifacts for public use after acceptance.
Dataset Considerations  We use publicly available datasets that underwent safety checks such as FineWeb-EDU.
Documentaion of Artifacts  Our work generates English text, and our codebase is primarily in Python.
@@ 8 Ethics Statement
Package Usage  we use nltk for BLEU, NIST, METEOR, and N gram diversity. We use https://pypi.org/project/codebleu/ for codebleu.

### v2

@@ 4.4 Training Objective
$\mathcal{L}_{\mathrm{pos}}(\theta)\;=\;\mathbb{E}_{t}\!\left[\,\big\|\,v_{\theta}(z_{t},x_{t},t)\;-\;\bigl(z_{0}-\tilde{z}_{T}\bigr)\big\|_{2}^{2}\right].$ (9)
Token loss.
For tokens we use the SEDD score-entropy objective from section 3, denoted $\mathcal{L}_{\mathrm{tok}}(\theta)$.
Total loss.
The final objective is a weighted sum
$\mathcal{L}(\theta)\;=\;\mathcal{L}_{\mathrm{tok}}(\theta)\;+\;\lambda\,\mathcal{L}_{\mathrm{pos}}(\theta),$ (10)
with $\lambda$ controlling the relative weight of position supervision.
Initialization: DDOT-R vs. DDOT-U.
During inference we manipulate how we sample $z_{T}$. By default (DDOT-R), we randomly sample a full length-$L$ terminal vector $z_{T}\sim\mathcal{U}(-1,1)^{L}$, reserve any $\ell^{p}$ entries as the prompt slice $z_{T}^{p}$. However, random sampling can produce high-density regions that tend to map to pad tokens because there are fewer tokens in the corresponding ground-truth region.
To mitigate this, DDOT-U places terminal positions on uniform grids for each set:
$z^{p}_{T,i}\;=\;\frac{\,2i-(\ell^{p}-1)\,}{\max\{1,\;\ell^{p}-1\}},$ $i=0,\dots,\ell^{p}-1,$ (11)
$z^{np}_{T,i}\;=\;\frac{\,2i-(\ell^{np}-1)\,}{\max\{1,\;\ell^{np}-1\}},$ $i=0,\dots,\ell^{np}-1.$
Uniform spacing avoids random clustering, reduces pad spillover, and yields more stable placements.
Reverse-time updates.
During sampling, positions follow near-straight trajectories because the learned velocity field approximates the constant ground-truth direction $z_{0}-\tilde{z_{T}}$. We therefore update positions with a simple Euler step at each $\tau$-leap of the SEDD reverse process, while keeping prompts clamped and pads stationary (Equation 5).
@@ 4.5 Simultaneous Text & Position Diffusion
DDOT performs discrete text and continuous position diffusion simultaneously, as these processes operate independently in continuous time. We therefore predict both token-value scores and position velocities in a single forward pass. This independence also enables simulation-free training by sampling token and position states independently at arbitrary time steps. We summarize the training procedure in algorithm 1.
Algorithm 1 DDOT Training
1:batch $(x_{0},z_{0})$; loss weight $\lambda$
2:Sample $t\sim\mathcal{U}(0,1)$
3:Build terminal positions: sample $z_{T}\sim\mathcal{U}(-1,1)^{L}$ and split $(z_{T}^{p},z_{T}^{\mathrm{np}})$, or use uniform grids from Equation 11
4:Compute within-set OT: $\pi^{p}$ via 1D sort-match (Equation 6); $\pi^{r}$ via injective matching (Equation 7); set $\tilde{z}_{T}^{p}=z_{T}^{p}[\pi^{p}]$, $\tilde{z}_{T}^{r}=z_{T}^{\mathrm{np}}[\pi^{r}]$, leftovers $\to z_{T}^{\mathrm{pad}}$
5:Form paths $z_{t}^{S}=(1-t)z_{0}^{S}+t\,\tilde{z}_{T}^{S}$ for $S\in\{p,r\}$; set pads via Equation 5
6:Sample token states $x_{t}$ from the SEDD forward at time $t$
7:Predict $s_{\theta}(x_{t},t)$ and $v_{\theta}(z_{t},x_{t},t)$
8:Compute $\mathcal{L}_{\mathrm{tok}}$ (Equation 2) and $\mathcal{L}_{\mathrm{pos}}$ (Equation 9)
9:Update $\theta$ on $\mathcal{L}_{\mathrm{tok}}+\lambda\,\mathcal{L}_{\mathrm{pos}}$ (Equation 10)
@@ 4.6 Implementation Details
We extend SEDD, which is based on the Diffusion Transformer architecture Peebles and Xie (2023), with two additional modules. First, we introduce a learnable type embedding applied directly after the token-embedding lookup. This embedding indicates whether a token is part of the prompt or the masked response ($x_{i}\in x^{p}$ or $x_{i}\in x^{n}p$), which is critical for assigning each token to the correct OT flow. Second, we add a linear head at the end of the Diffusion Transformer to compute $v_{\theta}(z_{t},t)$. The model architecture is available in Figure 2.
To incorporate continuous positional information, we scale $z_{t}$ from [1,1][-1,1] to match the context length of the original pretrained model (1024). We then use Rotary Position Embeddings Su et al. (2024), a standard technique in discrete diffusion models. Implementation details can be found in Subsection A.3.2.
@@ 5.1 Experimental Setup
Datasets   We evaluate our approach on the One-Billion-Word and Yelp datasets, following the preprocessing steps outlined in prior work on infilling and lexically constrained generation Miao et al. (2019); Zhang et al. (2020); Iso (2024). These datasets consist of examples with $1$ to $6$ keywords that must be infilled while maintaining their relative order to generate coherent sentences. In addition to randomly masking positions, we also introduce a block masking method that masks a single continuous span of text ranging from 0 to $L/2$ tokens (32 for One-Billion-Word and Yelp; 512 for CodeParrot). Finally, we apply the aforementioned masking methods to the Python subset of the CodeParrot dataset. Table 1 illustrates examples of this lexically constrained generation task.
@@
Model One Billion Word–Block Yelp-Block CodeParrot-Block CodeParrot-Random
@@
Model One Billion Word–Random Yelp-Random
@@
Training Details   To align the position-prediction modules, we first fine-tune SEDD with the added modules on FineWeb-Edu Penedo et al. (2024). Afterward, we further fine-tune on the One-Billion-Word and Yelp datasets. For simplicity, we keep all parameters unfrozen and optimize $\mathcal{L}_{\text{tok}}$ and $\mathcal{L}_{\text{pos}}$ simultaneously.
@@
In line with SEDD, we train our model in two configurations: small (90M non-embedding parameters) and medium (320M non-embedding parameters). DDOT-medium is on the same scale as CBART (406M parameters) and AutoTemplate-base (220M parameters). Following SEDD, we use the AdamW optimizer with a learning rate of $3\times 10^{-5}$. We set $\lambda=10$ for position-loss weighting. For each experiment, we use either $48\times$ L40 (48 GB), $80\times$ A30 (24 GB), or $8\times$ A100 (80 GB) GPUs.
@@
Baselines   We compare our method against strong autoregressive (AR) and non-autoregressive (NAR) baselines. AutoTemplate Iso (2024), a state-of-the-art AR approach, leverages the T5 Raffel et al. (2020) family of pretrained models and parses the lexically constrained generation task into a template that is generated autoregressively from left to right. The previous state-of-the-art NAR method, CBART He (2021), is built on the pretrained BART framework Lewis et al. (2020) and iteratively inserts tokens into a sequence.
@@
We also introduce two diffusion-based baselines that follow the same training procedure as DDOT. Left Context (LC) concatenates all prompt tokens to the left of the sequence and generates the response to the right of a separator token. Position Prediction (PoP) uses a SEDD model with a linear head that first predicts the positions of every token; this sequence is then fed through a fine-tuned fixed-position SEDD.
@@
Distribution Annealing   Many lexically constrained generation baselines, including AutoTemplate and CBART, use distribution-annealing methods such as top-$p$, top-$k$, greedy decoding, and beam search. To provide a parallel to greedy decoding—which always takes the top token probability—we anneal the token-value distribution during sampling to include only the most probable non-mask token. Specifically, given the predicted probability that a token is the mask, $\hat{p}(x^{\text{mask}})$, we assign $1-\hat{p}(x^{\text{mask}})$ to the token value with the highest probability excluding the mask token, and set the remaining token probabilities to 0. Greedy decoding in prior models (e.g., AR) is deterministic, collapsing the tree of all generation paths into a single path. However, our annealing process maintains generation diversity (A.2.2) because the model still samples from the annealed distribution over the top token value and the mask token. Whenever possible, we evaluate against the greedy-decoding baseline.
Metrics  Following prior work Miao et al. (2019); Zhang et al. (2020); Iso (2024), we evaluate BLEU-2/4 Papineni et al. (2002), NIST-2/4 Doddington (2002), METEOR-v1.5 Denkowski and Lavie (2014), and success rate (SR), which is the percentage of responses that are valid infilling results. A result is determined as invalid when any prompt tokens in the generated sequence do not appear in the same relative order that they were provided.
@@ 5.2 Main Results
We present lexically constrained generation results in Table 2. Our approach uses greedy annealing and is compared against greedy decoding wherever applicable, including the CBART greedy-decoding baseline. Our method achieves competitive performance with previous NAR models, approaching AR performance. Notably, our model achieves state-of-the-art performance on most metrics among the diffusion baselines. Our method performs well on block infilling, which may be more useful in real-world applications. Furthermore, we observe that DDOT scales well to longer sequences: it frequently generates valid responses that include all prompt words in the same relative order, as reflected in the SR. In contrast, diffusion baselines quickly generate invalid responses as the number of prompt tokens increases (Figure 3, Table 2). However, in benchmarks with six or fewer prompt tokens, diffusion baselines maintain high SR (Table 3). This may be because fixed-position models have room to correct generation when the ratio of prompt to response tokens is low.
Table 3 compares results with previous work in lexically constrained generation. Since pretrained diffusion models currently lag behind AR models—an issue not unique to DDOT—we focus on NAR models. DDOT performs on par with previous state-of-the-art models and achieves higher SR than all diffusion baselines. Although DDOT underperforms LC and PoP on some metrics, we argue that the One-Billion-Word–Random and Yelp–Random settings over-index on the unrealistic task of generating text from only 1166 randomly spaced tokens. Furthermore, Figure 3 shows the broader trend when DDOT is scaled to a larger number of prompt tokens.
@@ 5.3 Analysis
In this section, we investigate the effect of random versus uniform position initialization, the inclusion of OT coupling, and the impact of varying the number of sampling steps.
Figure 4: (a) and (b) show ground-truth token positions over time. Without OT (a), many line crossings indicate unstable permutations, whereas with OT coupling (b), trajectories are nearly straight throughout the denoising process. (c) Performance tends to increase with more sampling steps.
Position Initialization   In Table 2 and Table 3, we also explore the difference between DDOT-R and DDOT-U. In DDOT-R, pad tokens tend to cluster in high-density regions because OT finds no matches for them; in DDOT-U, pad tokens tend to be evenly spaced. We find that DDOT-U consistently outperforms DDOT-R.
OT Coupling   To demonstrate the importance of OT coupling between source and target positions, we retrain the small version of DDOT without OT coupling and provide a quantitative comparison in Table 4. Models trained with OT coupling consistently outperform those using independent (random) coupling. We hypothesize that OT coupling provides a stronger signal about token ordering throughout the diffusion process. Specifically, DDOT guarantees that the relative ordering of the prompt and generated tokens at any time step matches the original ordering. In contrast, independent coupling requires the model to infer the original ordering—a challenging task given the numerous plausible orderings that can result from interspersed prompt tokens.
@@
Position Over Time   We qualitatively compare ground-truth token paths during training in Figure 4. With OT coupling, token trajectories exhibit significantly fewer crossings, maintaining relative order throughout the generation process; in contrast, independent coupling frequently permutes tokens. Visualizations of token paths during inference are available in subsection A.4.
Number of Sampling Steps   One advantage of diffusion models over autoregressive models is the ability to trade compute for accuracy by varying the number of inference steps. Figure 4 shows how the number of sampling steps influences lexically constrained generation performance: as the number of sampling steps increases, performance improves.
Wall-Time Analysis   We evaluate the inference speed of DDOT against the diffusion baselines on the One-Billion-Word. Table 5 presents the wall-clock time per batch alongside BLEU-2 and BLEU-4 scores for increasing numbers of sampling steps.
DDOT demonstrates significantly better efficiency. For any given number of sampling steps, DDOT is not only faster than LC and competitive with PoP in raw speed, but also achieves substantially higher BLEU scores. Notably, LC must regenerate prompt tokens and therefore requires up to double the input sequence length. PoP also requires an additional forward pass to predict initial positions.
@@
The added modules that enable position prediction are lightweight, consisting of a linear head and two type embeddings. By contrast, the LC baseline requires double the context length of DDOT because it must regenerate prompt tokens.
The OT calculation is highly efficient, taking 16 minutes, 11 seconds on an Intel Xeon 8462Y+ 64-core processor for the 10-billion-token subset of FineWeb-Edu. In practice, we stream the dataset, caching several OT couplings in advance without needing to preprocess them. With caching, it takes 4 minutes, 30 seconds to run 1,000 training steps on an L40 GPU with a batch size of 256. Without caching, it takes 4 minutes, 27 seconds—a negligible difference.
@@ 6 Conclusion
In this work, we introduce DDOT, the first discrete diffusion model capable of flexible-length text infilling by jointly denoising token values and positions. By incorporating optimal-transport coupling, DDOT preserves token order while enabling dynamic position adjustment, addressing limitations in existing text diffusion models. Our experiments show that DDOT outperforms diffusion baselines and matches state-of-the-art non-autoregressive models.
Limitations
DDOT inherits many drawbacks from previous discrete text diffusion methods Lou et al. (2024); Austin et al. (2021); Sahoo et al. (2024). First, diffusion backbones still lag behind autoregressive backbones. Since diffusion for text is an emerging field, this performance gap is expected. Similarly, the fine-tuned versions of these models (e.g., DDOT vs. AutoTemplate) exhibit the same gap. Second, unlike autoregressive models, existing diffusion models cannot use a KV cache to store previous activations. DDOT also inherits risks common to large-scale text models.
Ethical Considerations
Artifacts The artifacts we used have public-use licenses. Furthermore, we plan to release our code artifacts for public use after acceptance.
Dataset Considerations We use publicly available datasets that underwent safety checks, such as FineWeb-Edu.
Documentation of Artifacts Our work generates English text, and our codebase is primarily in Python.
@@
Package Usage We use NLTK for BLEU, NIST, METEOR, and $n$-gram diversity. We use https://pypi.org/project/codebleu/ for CodeBLEU.
Acknowledgements
We thank Virginia Tech’s Advanced Research Computing (ARC) for computational resources and technical support, and the anonymous reviewers for their constructive feedback.